# Julian Martinez-Iriarte, and Gabriel Montes-Rojas

### v3

@@ A.5 Proof of Proposition 1
$o_{p}(1)$ $=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\begin{bmatrix}Z_{i}\left(D_{i}-Z_{i}^{\prime}\pi_{1}-X_{i}^{\prime}\pi_{2}\right)\\ X_{i}\left(D_{i}-Z_{i}^{\prime}\pi_{1}-X_{i}^{\prime}\pi_{2}\right)\\ \nabla_{\gamma}h(X_{i},Z_{i};\gamma)\left(V_{i}^{2}-1\right)h(X_{i},Z_{i};\gamma)^{3}\end{bmatrix}$
$+\begin{bmatrix}-E[Z_{i}Z_{i}^{\prime}]&-E[Z_{i}X_{i}^{\prime}]&0\\ -E[X_{i}Z_{i}^{\prime}]&-E[X_{i}X_{i}^{\prime}]&0\\ 0&0&-2E[\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\nabla_{\gamma}h(X_{i},Z_{i};\gamma)^{\prime}h(X_{i},Z_{i};\gamma)^{2}]\\ \end{bmatrix}$
$\sqrt{n}\begin{bmatrix}\widehat{\pi}_{1}-\pi_{1}\\ \widehat{\pi}_{2}-\pi_{2}\\ \widehat{\gamma}-\gamma\end{bmatrix}.$
Therefore, the influence function of $\sqrt{n}(\widehat{\phi}-\phi)$ is
$\sqrt{n}(\widehat{\phi}-\phi)$ $=\begin{bmatrix}E[Z_{i}Z_{i}^{\prime}]&E[Z_{i}X_{i}^{\prime}]&0\\ E[X_{i}Z_{i}^{\prime}]&E[X_{i}X_{i}^{\prime}]&0\\ 0&0&2E[\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\nabla_{\gamma}h(X_{i},Z_{i};\gamma)^{\prime}h(X_{i},Z_{i};\gamma)^{2}]\\ \end{bmatrix}^{-1}$
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\begin{bmatrix}Z_{i}\left(D_{i}-Z_{i}^{\prime}\pi_{1}-X_{i}^{\prime}\pi_{2}\right)\\ X_{i}\left(D_{i}-Z_{i}^{\prime}\pi_{1}-X_{i}^{\prime}\pi_{2}\right)\\ \nabla_{\gamma}h(X_{i},Z_{i};\gamma)\left(V_{i}^{2}-1\right)h(X_{i},Z_{i};\gamma)^{3}\end{bmatrix}+o_{p}(1).$
Since the matrix above is block diagonal, then we have
$\sqrt{n}(\widehat{\phi}-\phi)$ $=\begin{bmatrix}\begin{bmatrix}E[Z_{i}Z_{i}^{\prime}]&E[Z_{i}X_{i}^{\prime}]\\ E[X_{i}Z_{i}^{\prime}]&E[X_{i}X_{i}^{\prime}]\end{bmatrix}^{-1}&0\\ 0&\frac{1}{2}E[\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\nabla_{\gamma}h(X_{i},Z_{i};\gamma)^{\prime}h(X_{i},Z_{i};\gamma)^{2}]^{-1}\\ \end{bmatrix}$
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\begin{bmatrix}Z_{i}\left(D_{i}-Z_{i}^{\prime}\pi_{1}-X_{i}^{\prime}\pi_{2}\right)\\ X_{i}\left(D_{i}-Z_{i}^{\prime}\pi_{1}-X_{i}^{\prime}\pi_{2}\right)\\ \nabla_{\gamma}h(X_{i},Z_{i};\gamma)\left(V_{i}^{2}-1\right)h(X_{i},Z_{i};\gamma)^{3}\end{bmatrix}+o_{p}(1).$
The inverses of the blocks exists because of Assumptions 2 and 7. Asymptotic normality follows from Assumption 7 by the second moments being finite. ∎
@@ A.6 Proof of Theorem 1
where $U:=Y-E[Y|D,W]$ and satisfies, by construction, $E[U|D,W]=0$. We follow Appendix 6A in Wooldridge (2010) closely. To implement the regression, $V$ is replaced by $\widehat{V}$. Recall that $V$ is given by
@@
If $\phi$ were known, then the influence function of $\alpha$ would be given by $\Sigma_{\alpha}^{-1}R_{i}(\phi)U_{i}$ where $\Sigma_{\alpha}:=E[R_{i}(\phi)R_{i}(\phi)^{\prime}]$. To account for the influence of estimating $\phi$ we need to evaluate the effect is has on $R(\phi)U=R(\phi)(Y-R(\phi)^{\prime}\alpha)$.
@@
$\frac{1}{n}\sum_{i=1}^{n}R_{i}(\widehat{\phi})(Y_{i}-R_{i}(\widehat{\phi})^{\prime}\widehat{\alpha})$ $=\frac{1}{n}\sum_{i=1}^{n}R_{i}(\phi)(Y_{i}-R_{i}(\phi)^{\prime}\widehat{\alpha})$
@@
$=\frac{1}{n}\sum_{i=1}^{n}R_{i}(\phi)(Y_{i}-R_{i}(\phi)^{\prime}\widehat{\alpha})+E[UJ_{\phi}-R(\phi)\alpha^{\prime}J_{\phi}](\widehat{\phi}-\phi)+o_{p}(n^{-1/2}),$
$=\Sigma_{\alpha}(\widehat{\alpha}-\alpha)+E[UJ_{\phi}-R(\phi)\alpha^{\prime}J_{\phi}](\widehat{\phi}-\phi)+o_{p}(n^{-1/2}).$
@@
$\psi_{i}=R_{i}U_{i}+E[UJ_{\phi}-R(\phi)\alpha^{\prime}J_{\phi}]\Sigma_{\phi}^{-1}M_{i}$
and $\Psi:=E[\psi\psi^{\prime}]$. Therefore,
$\sqrt{n}(\widehat{\alpha}-\alpha)$ $=\Sigma_{\alpha}^{-1}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\psi_{i}+o_{p}(1),$

### v4

@@ A.5 Proof of Proposition 1
$E\left[\frac{\partial m_{i}(\phi)}{\partial\phi^{\prime}}\right]=\begin{bmatrix}-E[Z_{i}Z_{i}^{\prime}]&-E[Z_{i}X_{i}^{\prime}]&0\\ -E[X_{i}Z_{i}^{\prime}]&-E[X_{i}X_{i}^{\prime}]&0\\ 0&0&-2E[\nabla_{\gamma}h_{i}\nabla_{\gamma}h_{i}^{\prime}h_{i}^{2}]\\ \end{bmatrix}$.
Therefore, a LLN holds for $\frac{\partial m_{i}(\phi)}{\partial\phi^{\prime}}$ and $E[\frac{\partial m_{i}(\phi)}{\partial\phi^{\prime}}]$ is invertible because of Assumptions 2 and 7. Therefore, we have
$0$ $=\frac{1}{n}\sum_{i=1}^{n}m_{i}(\phi)+E\left[\frac{\partial m_{i}(\phi)}{\partial\phi^{\prime}}\right](\widehat{\phi}-\phi)+o_{p}(\|\widehat{\phi}-\phi\|)$
$=\frac{1}{n}\sum_{i=1}^{n}m_{i}(\phi)-\Sigma_{\phi}(\widehat{\phi}-\phi)+o_{p}(\|\widehat{\phi}-\phi\|),$
where $\Sigma_{\phi}:=-E\left[\frac{\partial m_{i}(\phi)}{\partial\phi^{\prime}}\right]$. Multiplying through by $\sqrt{n}$ and solving for $\widehat{\phi}-\phi$ yields
$\sqrt{n}(\widehat{\phi}-\phi)$ $=\Sigma_{\phi}^{-1}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}m_{i}(\phi)+o_{p}(1).$
Since the $M$ matrix above is block diagonal, then we have
$\Sigma_{\phi}^{-1}=\begin{bmatrix}\begin{bmatrix}E[Z_{i}Z_{i}^{\prime}]&E[Z_{i}X_{i}^{\prime}]\\ E[X_{i}Z_{i}^{\prime}]&E[X_{i}X_{i}^{\prime}]\end{bmatrix}^{-1}&0\\ 0&\frac{1}{2}E[\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\nabla_{\gamma}h(X_{i},Z_{i};\gamma)^{\prime}h(X_{i},Z_{i};\gamma)^{2}]^{-1}\end{bmatrix}$.
Then, $\psi_{\phi,i}=\Sigma_{\phi}^{-1}m_{i}$, and $\sqrt{n}(\widehat{\phi}-\phi)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\psi_{\phi,i}+o_{p}(1).$ Therefore, $\sqrt{n}(\widehat{\phi}-\phi)\overset{d}{\to}\mathcal{N}(0,E[\psi_{\phi}\psi_{\phi}^{\prime}]).$ Thus, $\Omega_{\phi}=E[\psi_{\phi}\psi_{\phi}^{\prime}].$
@@ A.6 Proof of Theorem 1
where U:=YE[Y|D,W]U:=Y-E[Y|D,W] and is implemented with $V$ replaced by $\widehat{V}$. We follow Appendix 6A in Wooldridge10 closely. Recall that $V$ is given by
@@
If $\phi$ were known, then the influence function of $\hat{\alpha}$ would be given by $E[R(\phi)R(\phi)^{\prime}]^{-1}R_{i}(\phi)U_{i}$. To account for the influence of estimating $\phi$ we need to evaluate the effect is has on $R(\phi)U=R(\phi)(Y-R(\phi)^{\prime}\alpha)$.
@@
[…]
@@
$=\frac{1}{n}\sum_{i=1}^{n}R_{i}(\phi)(Y_{i}-R_{i}(\phi)^{\prime}\widehat{\alpha})+E[UJ_{\phi}-R(\phi)\alpha^{\prime}J_{\phi}](\widehat{\phi}-\phi)+o_{p}(n^{-1/2})$
$=\frac{1}{n}\sum_{i=1}^{n}R_{i}(\phi)U_{i}-\left(\frac{1}{n}\sum_{i=1}^{n}R_{i}(\phi)R_{i}(\phi)^{\prime}\right)(\widehat{\alpha}-\alpha)$
$\ \ +E\left[UJ_{\phi}-R(\phi)\alpha^{\prime}J_{\phi}\right](\widehat{\phi}-\phi)+o_{p}(n^{-1/2}).$
@@
$\psi_{\alpha,i}=E[RR^{\prime}]^{-1}R_{i}U_{i}+E[RR^{\prime}]^{-1}E[UJ_{\phi}-R\alpha^{\prime}J_{\phi}]\psi_{\phi,i}$
and $\Omega_{\alpha}:=E[\psi_{\alpha}\psi_{\alpha}^{\prime}]$. Therefore,
$\sqrt{n}(\widehat{\alpha}-\alpha)$ $=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\psi_{\alpha,i}+o_{p}(1),$